# Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Writing - original draft, Writing - review & editing, Visualization

### v2

@@ Abstract
Multiple instance learning (MIL) has emerged as a powerful framework for weakly supervised whole slide image (WSI) classification, enabling slide-level predictions without requiring detailed patch-level annotations. However, a key limitation of MIL lies in the underexplored potential of pre-training the MIL aggregator. Most existing approaches train it from scratch, resulting in performance heavily dependent on the number of labeled WSIs, while overlooking the abundance of unlabeled WSIs available in real-world scenarios. To address this, we propose PreMix, a novel framework that leverages a non-contrastive pre-training method, Barlow Twins, augmented with the Slide Mixing approach to generate additional positive pairs and enhance feature learning, particularly under limited labeled WSI conditions. Fine-tuning with Mixup and Manifold Mixup further enhances robustness by effectively handling the diverse sizes of gigapixel WSIs. Experimental results demonstrate that integrating HIPT into PreMix achieves an average F1 improvement of 4.7% over the baseline HIPT across various WSI training datasets and label sizes. These findings underscore its potential to advance WSI classification with limited labeled data and its applicability to real-world histopathology practices. The code is available at https://anonymous.4open.science/r/PreMix
@@
{highlights}
This study introduces PreMix, a framework that leverages the Barlow Twins method with a Slide Mixing approach to pre-train MIL aggregators, addressing the scarce labeled WSIs and the underexplored potential of utilizing unlabeled WSIs in weakly supervised classification.
Fine-tuning with Mixup and Manifold Mixup further enhances robustness by effectively handling the diverse sizes of gigapixel WSIs.
Experimental results demonstrate that integrating HIPT into PreMix achieves a 4.7% mean F1 improvement over the baseline HIPT, showcasing its scalability and effectiveness across varying WSI datasets and labeling scenarios.
@@ 1 Introduction
Histopathology, the study of tissues under a microscope, is essential for diagnosing diseases, predicting outcomes, and planning treatments [1]. Traditionally, this process relied heavily on the expertise of trained professionals analyzing tissue samples. However, advancements in medical imaging have significantly transformed the field. One of the most notable developments is the use of whole slide images (WSIs), which digitally capture detailed views of tissue samples for digital pathology [2]. These high-resolution images offer deep insights into tissue structures and abnormalities, enabling more effective diagnosis and treatment planning [3, 4]. However, this advancement comes with complexities. The vast scale of WSIs, characterized by gigapixel resolution, and the often subtle disease indications pose challenges for both computer professionals and medical practitioners [5].
The immense size of WSIs makes it impractical to directly train deep-learning models on them without preprocessing into smaller patches, given current GPU limitations. Additionally, annotating WSIs is time-consuming and costly due to their resolution, and manual labeling is prone to errors and inconsistencies [6]. To address these challenges, multiple instance learning (MIL) offers a weakly-supervised approach for WSI classification, particularly useful for clinical pre-screening. MIL accommodates varying WSI sizes and relies only on slide-level labels, eliminating the need for labor-intensive patch-level annotations. By reducing dependency on detailed labels, MIL overcomes the limitations of traditional supervised methods and provides a more robust and efficient solution for WSI analysis [7].
@@
As shown in Figure 1(a), the MIL process typically involves tiling WSIs into patches, processing these patches through a pre-trained feature extractor (e.g., CNN or ViT), and generating either scalar outputs (instance-based MIL) or feature vectors (embedding-based MIL). Among these approaches, embedding-based MIL, which provides richer representations, has been shown to be more effective for WSI classification [8]. The outputs are then aggregated using an MIL aggregator for slide-level prediction. However, existing MIL approaches [9, 10, 11, 12, 13, 14, 15] typically train the MIL aggregator from scratch, which limits performance when only a small number of labeled WSIs are available.
We argue that improving WSI classification performance with a limited number of labeled training WSIs requires leveraging the abundance of unlabeled WSIs, reflecting the common real-world scenario where many WSIs remain unlabeled. By utilizing these unlabeled WSIs and employing a proper pre-training method for the MIL aggregator, we aim to achieve better initialization, leading to enhanced WSI classification performance. While we are not the first to explore MIL aggregator pre-training, SSMIL [16] pioneered the use of the SimCLR [17] contrastive method for pre-training MIL aggregators at the WSI level. However, this approach has been shown to underperform compared to training the MIL aggregator from scratch. We attribute this limitation to the inherent imbalance in gigapixel WSI datasets (e.g., a significantly larger number of normal WSIs compared to tumor WSIs), which often leads to the inclusion of incorrect or noisy negative pairs—a critical issue for contrastive methods reliant on both positive and negative pairs.
To address these challenges, we propose the PreMix (Pre-training and Mixing) framework (see Figure 1(b)). The core innovation of PreMix lies in adopting a non-contrastive approach, Barlow Twins [18], as the base pre-training method. Barlow Twins relies solely on positive pairs, eliminating the need for negative pairs. We extend this method with an intra-batch mixing strategy, termed Barlow Twins Slide Mixing, to generate additional positive pairs from unlabeled WSIs. This approach leverages the semantic relatedness of mixed slide features, even when they originate from different slides within the same batch, to enrich the pre-training process.
To further enhance robustness and generalization, we incorporate Mixup [19] and Manifold Mixup [20] during the fine-tuning stage. These techniques are adapted to handle the unique challenges of WSIs, enabling the mixing of features and labels across slides of varying sizes within the batch. The PreMix framework is evaluated using HIPT [21] as the baseline, and we demonstrate its effectiveness by integrating HIPT into PreMix to validate its generalizability across different training datasets and sample sizes.
@@
  • We identify the limitations of contrastive learning methods (e.g., SimCLR) for MIL aggregator pre-training, as they hinder WSI classification performance. Instead, we propose the use of non-contrastive learning methods (e.g., Barlow Twins), which do not require negative pairs and are more suitable for imbalanced gigapixel WSI datasets, achieving better classification results.
  • We introduce PreMix, an extension of the general MIL framework, which enhances the MIL aggregator by pre-training it with the novel Barlow Twins Slide Mixing approach. This method extends Barlow Twins by incorporating intra-batch mixing to generate additional positive pairs, improving feature representation during pre-training.
  • We integrate Mixup and Manifold Mixup during MIL aggregator fine-tuning, enabling WSI feature and label mixing across different WSI feature sizes. Furthermore, we demonstrate that applying Manifold Mixup at the Transformer Encoder layer yields the most significant performance improvements.
  • To validate the proposed MIL aggregator pre-training method under limited labeling scenarios, we conduct experiments with five different WSI training sample sizes. Additionally, we evaluate its robustness across datasets using random sampling (traditional fully-supervised fine-tuning) and five active learning acquisition functions: Entropy, K-means++, Coreset, BADGE, and CDAL. These experiments demonstrate the framework’s robustness across varying WSI datasets and training sizes, highlighting its versatility and effectiveness.
The remainder of this paper is organized as follows: Section 2 introduces related literature, including self-supervised learning for pre-training in the histopathology domain, MIL in histopathology domain at both single and multi scales, existing data augmentation strategies. Section 3 details the problem formulation, overall training process, and the proposed framework. Section 4 describes the dataset, implementation setup, model evaluation, experiment results, and ablation studies. In Section 5, we discuss our findings and future research directions. Finally, we conclude our paper in Section 6.
@@ 2.1 Self-supervised learning in histopathology
In the medical field, where obtaining expert annotations is challenging and expensive, self-supervised learning (SSL) has emerged as a promising solution since it does not rely on labeled data [22]. Ciga et al. [23] demonstrated the effectiveness of the SimCLR method [17] for pre-training models in histopathology. Their work highlighted the robustness of patch-level feature representations for various downstream tasks. Building on this, Chen and Krishnan [24] proposed replacing SimCLR-trained ResNet-50 encoders [25] with ViTs [26], leveraging the DINO method [27] to better capture intricate details of histopathological tissues. Kang et al. [28] further conducted a comprehensive evaluation of SSL methods in histopathology, comparing four representative SSL approaches: Barlow Twins [18], SwAV [29], MoCo V2 [30], and DINO [27]. Recently, Wong and Yi [31] provided practical guidance on optimizing WSI classification by analyzing feature extractors from three dimensions: pre-training dataset, backbone model, and pre-training method.
While most SSL studies in histopathology focus on learning patch-level feature representations, SSMIL [16] stands out as the first to explore pre-training the MIL aggregator using SimCLR [17]. However, MIL aggregators fine-tuned using this approach still underperform compared to those trained from scratch, highlighting the need for more effective pre-training strategies.
@@ 2.2 Multiple instance learning for WSI analysis
Recent advancements in multiple instance learning (MIL) for WSI classification have moved beyond traditional aggregation methods, such as mean or max pooling, to techniques that extract and combine patch-level features for slide-level predictions [32]. Early methods like MIL-RNN [9] aggregates patch features across slides but struggled with long-distance information loss due to the large number of patches. To address these spatial complexities, ABMIL [10] computes attention scores for each patch, while DSMIL [11] applies distances between instances and a critical instance as weights. CLAM [12] refines feature spaces through instance-level clustering, while TransMIL [33] uses a transformer architecture to capture relationships among patches. DTFD-MIL [34] addresses data limitations in WSIs by introducing pseudo-bags and calculating instance probabilities under the ABMIL framework. Recently, MHIM-MIL [35] focuses on identifying hard-to-classify instances using a momentum teacher to train attention-based MIL models.
Despite their effectiveness, single-scale methods are limited by their inability to capture spatial correlations across resolutions. Multi-resolution strategies address this limitation by leveraging features from multiple magnifications. DSMIL-LC [11] concatenates low- and high-resolution features to improve context-awareness. HIPT [21] employs a hierarchical transformer to extract detailed and contextual features across scales, while MS-DA-MIL [36], H2-MIL [37], and DAS-MIL [38] use graph-based models to enhance spatial and scaling relationships between patches.
HIPT’s hierarchical feature extraction process stands out for its ability to efficiently combine detail and context across scales, making it particularly suitable for both pre-training and downstream classification tasks. Motivated by these advantages, we adopt the original HIPT as the baseline and evaluate its performance within the PreMix framework to demonstrate its effectiveness.
@@ 2.3 Data augmentation
To address the challenges of limited data and improve deep learning performance, various data augmentation techniques have been developed [39]. One prominent method is Mixup [19], which blends images and their labels to enhance classification, inspiring numerous domain-specific techniques such as Balanced Mixup [40] and Stain Mixup [41]. The integration of such data mixing methods into SSL has led to significant improvements. For example, Ren et al. [42] demonstrated that incorporating mixed natural images as supplementary positive pairs during pre-training enhanced the performance of MoCo V2.
While many augmentation techniques are designed for natural images or specific medical applications, the unique challenges of WSIs, such as their gigapixel size and diverse feature scales, have led to the development of specialized solutions. Methods like ReMix [32], RankMix [43], and PseMix [44] focus on downstream classification, adapting data augmentation techniques to the WSI domain. In contrast, our approach incorporates data augmentation during both pre-training and downstream classification. Specifically, we employ an intra-batch mixing approach during the pre-training of the MIL aggregator at the WSI level, enabling the generation of additional positive pairs while handling diverse WSI feature sizes. In downstream classification, we adapt Mixup [19] and Manifold Mixup [20] to combine WSI features and labels of different sizes. Originally developed for natural images of consistent size, these methods are modified to address the unique characteristics of WSIs, improving performance and robustness.
@@ 3.1 General MIL framework
Figure 2: The illustration of our training process utilizing the HIPT feature extraction method showcases the selection of WSI candidates for labeling using either random sampling (traditional fully-supervised fine-tuning) or active learning. The selected WSIs are weakly-labeled by pathologists and incorporated into the next training iteration. A key distinction between the baseline HIPT and our HIPT with PreMix is that the latter incorporates MIL aggregator initialization from Barlow Twins Slide Mixing pre-training and employs Mixup and Manifold Mixup during fine-tuning, while the baseline does not.
@@ 3.1.2 Feature aggregation
After extraction, the individual patch features are aggregated to form a slide-level representation. This feature aggregation process is facilitated by the MIL aggregator, $f_{a}$. In most cases, $f_{a}$ is trained from scratch. This process yields a comprehensive slide-level representation, expressed as $f_{a}(P_{i})$in $\mathbb{R}^{1\times d}$. The primary objective of this step is to conduct the binary classification of slides. Consequently, the final representations are fed into a linear classifier to distinguish between two classes. Each slide is assigned a label $Y_{i}\in\{0,1\}$, where a slide is labeled as $Y_{i}=0$ if it is normal and $Y_{i}=1$ if it contains tumor.
@@ 3.2 HIPT feature extraction
Figure 3: Barlow Twins Slide Mixing: This process relies on region features extracted via HIPT as inputs, accommodating varying WSI input sizes through zero-padding and masking techniques and the addition of intra-batch mixing strategy during pre-training. "Flip" means from the same batch but with the batch order reversed.
@@ 3.3 Overall training process
Finally, the most informative WSIs, determined by active learning acquisition functions or random sampling are selected from the unlabeled pool. These selected WSIs are then labeled by pathologists in a weakly-supervised setting, requiring only slide-level labels without fine-grained annotations. Subsequently, they are added to the training data for the next cycle. This process continues until the total labeling budget for WSIs (number of available WSIs and their labels) is exhausted.

### v3

@@ Abstract
Multiple instance learning (MIL) has emerged as a powerful framework for weakly supervised whole slide image (WSI) classification, enabling slide-level predictions without requiring detailed patch-level annotations. Despite its success, a critical limitation of current MIL methods lies in the underutilization of pre-training for the MIL aggregator. Most existing approaches initialize the aggregator randomly and train it from scratch, making performance highly sensitive to the quantity of labeled WSIs and ignoring the abundance of unlabeled WSIs commonly available in clinical settings. To address this, we propose PreMix, a novel framework that leverages a non-contrastive pre-training method, Barlow Twins, augmented with the Slide Mixing approach to generate additional positive pairs and enhance feature learning, particularly under limited labeled WSI conditions. Fine-tuning with Mixup and Manifold Mixup further enhances robustness by effectively handling the diverse sizes of gigapixel WSIs. Experimental results demonstrate that integrating PreMix as a plug-in module into HIPT yields an average F1 improvement of 4.7% over the baseline HIPT across various WSI training sizes and datasets. These findings underscore its potential to advance WSI classification with limited labeled data and its applicability to real-world histopathology practices. The code is available at https://github.com/bryanwong17/PreMix
@@ 1 Introduction
Histopathology provides microscopic insights into tissue structures, making it essential for diagnosing diseases, predicting outcomes, and guiding treatment [1]. Traditionally, this work has relied on the expertise of pathologists who examine tissue samples under a microscope, a process that can be time-consuming and subject to interpretation. In recent years, digital pathology has emerged as a major advancement, with whole slide images (WSIs) offering high-resolution digital scans of tissue slides [2]. These images provide a comprehensive view of tissue architecture and cellular details, helping clinicians make more accurate and consistent decisions [3, 4].
Despite their potential, WSIs introduce new challenges. Their extremely large size, often reaching gigapixel resolution, makes it difficult to train deep learning models directly without dividing them into smaller patches due to hardware limitations. Moreover, annotating these images requires significant time and effort from medical professionals, and manual labeling can lead to inconsistencies across cases [5]. To address these issues, multiple instance learning (MIL) has gained attention as a weakly supervised approach for WSI classification. MIL enables models to learn from slide-level labels without needing detailed annotations for each patch, making it well-suited for clinical workflows where labeling resources are limited [6].
@@
As shown in Figure 1(a), the MIL pipeline typically begins by dividing WSIs into smaller patches, which are then processed by a pre-trained feature extractor such as a CNN [7] or ViT [8]. This results in either scalar scores in instance-based MIL or feature embeddings in embedding-based MIL, with the latter demonstrating better performance in WSI classification due to its ability to capture richer and more expressive representations [9]. These embeddings are subsequently combined using an MIL aggregator to generate slide-level predictions.
Although prior work has focused extensively on improving the quality of patch-level feature extractors [10, 11, 12, 13], most MIL models [14, 15, 16, 17, 18, 19, 20, 21, 22, 23] still train the aggregator from random initialization. As a result, their performance tends to rely heavily on the availability of large, labeled WSI datasets, which are often scarce in real-world clinical applications due to privacy constraints [24], the rarity of certain diseases [25, 26], and the high cost of obtaining weak labels from domain experts [27, 28]. On average, an experienced pathologist needs over 20 to 30 minutes to examine and assign weak labels to a single slide [29], making large-scale curation highly resource intensive.
To address these limitations, a few studies have explored pre-training strategies for MIL aggregators. For example, SSMIL [30] applies a self-supervised contrastive learning approach using SimCLR [31] at the WSI level. However, these methods have not consistently outperformed training from scratch. One possible reason is the inherent class imbalance in most WSI datasets, where negative slides (e.g., normal tissue) are much more prevalent than positive slides with tumors. This imbalance raises the risk of introducing noisy or misleading negative pairs, posing a major challenge for contrastive methods that depend on balanced positive and negative examples to learn meaningful representations.
As an alternative to contrastive learning methods that rely on negative pairs, we propose PreMix (Pretraining and Mixing), a self-supervised learning (SSL) framework at the WSI level designed to reduce sensitivity to class imbalance. PreMix adopts Barlow Twins [32], a non-contrastive method that learns representations using only positive pairs, thereby avoiding issues related to negative sample selection. To further improve the effectiveness of pre-training, we introduce Barlow Twins Slide Mixing, an intra-batch mixing strategy that interpolates features across slides within the same batch to create new positive pairs. This encourages semantic diversity and enhances representation quality, even in the absence of labeled data.
To complement the PreMix pre-training stage, we further enhance performance during fine-tuning by incorporating Mixup [33] and Manifold Mixup [34], which are adapted to the unique challenges of gigapixel WSIs. These mixing strategies allow for effective interpolation of both features and labels across slides of varying sizes, helping the MIL aggregator generalize better under limited supervision. By integrating both unsupervised and supervised mixing strategies, PreMix provides a unified framework that strengthens representation learning throughout the training pipeline. We validate its effectiveness by applying PreMix to HIPT [35], demonstrating consistent performance gains across training datasets and annotation budgets.
@@
  • We identify the limitations of contrastive learning (e.g., SimCLR) for MIL aggregator pre-training, especially in imbalanced WSI datasets, and instead adopt non-contrastive learning (Barlow Twins), which avoids reliance on negative pairs and yields better classification performance.
  • We propose PreMix, a novel MIL framework that introduces Barlow Twins Slide Mixing, an intra-batch feature interpolation strategy that generates diverse positive pairs and enhances representation learning during pre-training.
  • To further improve downstream performance, we integrate Mixup and Manifold Mixup during MIL aggregator fine-tuning, enabling feature and label mixing across slides of varying sizes. The Transformer encoder layer is found to yield the most significant performance gains.
  • We validate PreMix across five WSI training sizes and datasets using both random sampling (traditional fully supervised fine-tuning) and five active learning strategies: Entropy [36], K-means++ [37], Coreset [38], BADGE [39], and CDAL [40], demonstrating strong robustness under limited supervision.
@@ 2.1 Self-supervised learning in histopathology
In the medical domain, where expert annotations are often scarce and expensive, SSL has emerged as a powerful alternative to fully supervised approaches [41]. In histopathology, Ciga et al. [42] showed that models pre-trained with SimCLR [31] can extract strong patch-level features that generalize well across various downstream tasks. Building on this idea, Chen and Krishnan [43] introduced ViTs [8] into the pipeline, replacing ResNet-50 encoders [7] with DINO-trained backbones [44] to better capture complex tissue structures. Kang et al. [45] expanded this investigation by benchmarking several SSL methods, including Barlow Twins [32], SwAV [46], MoCo V2 [47], and DINO [44], across diverse histopathology tasks. More recently, Wong et al. [13] provided a practical perspective by examining how the choice of pretraining dataset, backbone architecture, and learning objective affects performance in WSI classification.
While most SSL studies in histopathology have focused on learning patch-level representations, less attention has been paid to the pretraining of MIL aggregators. One notable exception is SSMIL [30], which applied SimCLR to pre-train the MIL aggregator using WSI-level inputs. However, the performance of this contrastive approach remains inferior to models trained from scratch, suggesting that standard contrastive losses may not be well suited for the class imbalance commonly found in WSI datasets. This limitation highlights the need for alternative pretraining strategies that are more robust and compatible with the MIL setting.
@@ 2.2 Multiple instance learning for WSI analysis
Recent progress in MIL for WSI classification has shifted from simple aggregation techniques, such as mean or max pooling, to more sophisticated approaches that learn to combine patch-level features for slide-level predictions [48]. Early efforts like MIL-RNN [14] attempted to aggregate sequential patch features but were limited by the inability to capture long-range dependencies, especially in slides with a large number of patches. To address this, attention-based methods have been proposed. ABMIL [15] assigns attention weights to individual patches, while DSMIL [16] computes instance-level scores based on their distances to a representative “critical instance.” CLAM [17] enhances the feature space through instance-level clustering, and TransMIL [49] employs transformer-based architectures to model global relationships among patches. DTFD-MIL [50] tackles label sparsity by generating pseudo-bags and estimating instance probabilities within the ABMIL framework. More recently, MHIM-MIL [51] identifies challenging instances through a momentum teacher to refine attention-based MIL training.
Although effective, these methods often operate at a single resolution and may fail to capture multi-scale spatial dependencies inherent in WSIs. To address this limitation, multi-resolution MIL strategies have been introduced. DSMIL-LC [16] improves context-awareness by concatenating features from low and high magnifications. HIPT [35] advances this idea by employing a hierarchical transformer that extracts fine-grained and contextual representations across multiple resolutions. Similarly, models such as MS-DA-MIL [52], H2-MIL [53], and DAS-MIL [54] leverage graph-based architectures to model both spatial and scale-aware relationships among patches. More recently, HiVE-MIL [55] extended this line of work by integrating vision-language models (VLMs) into a hierarchical graph structure, jointly capturing intra- and inter-scale multimodal interactions between visual and textual modalities.
Among existing methods, HIPT stands out by using a hierarchical feature extractor that integrates information across different magnifications, effectively capturing both local details and global context through efficient feature representations. This makes HIPT particularly appealing for pretraining, especially in resource-constrained settings where GPU capacity is limited. Motivated by this, we use the HIPT feature extractor during pretraining and adopt the full HIPT architecture as our baseline to assess how well the proposed PreMix framework performs when applied as a plug-in module.
@@ 2.3 Data augmentation
To overcome the limitations of small datasets and improve model generalization, a wide range of data augmentation strategies have been proposed in deep learning [56]. Among them, Mixup [33] is a well-known method that mixes both inputs and their labels to create interpolated training examples, which has inspired several domain-specific variants such as Balanced Mixup [57] and Stain Mixup [58]. These techniques have also been extended to SSL. For instance, Ren et al. [59] showed that using mixed natural images as supplementary positive pairs during pretraining can significantly improve the performance of contrastive frameworks like MoCo V2 [47].
While most of these augmentation methods are developed for natural images or specific medical imaging tasks, WSIs present unique challenges due to their gigapixel scale and highly variable spatial resolution. To address this, several augmentation strategies tailored for WSIs have been introduced [48, 60, 61]. However, these methods are typically applied only during supervised learning.
In contrast, our approach integrates data mixing at both the pretraining and fine-tuning stages. During MIL aggregator pretraining, we introduce an intra-batch mixing strategy at the WSI level to generate additional positive pairs, helping the model better capture intra-cohort variability. For downstream classification, we adapt Mixup [33] and Manifold Mixup [34] to handle variable-sized WSI features and labels. Although originally developed for fixed-size natural images, we modify these techniques to suit the scale and diversity of WSIs, resulting in improved performance and robustness under limited supervision.
@@ 3.1 General MIL framework
Figure 2: The illustration of our training process utilizing the HIPT feature extraction method showcases the selection of WSI candidates for labeling using either random sampling (traditional fully supervised fine-tuning) or active learning. The selected WSIs are weakly-labeled by pathologists and incorporated into the next training iteration. A key distinction between the baseline HIPT and our HIPT with PreMix is that the latter incorporates MIL aggregator initialization from Barlow Twins Slide Mixing pre-training and employs Mixup and Manifold Mixup during fine-tuning, while the baseline does not.
@@ 3.1.2 Feature aggregation
After extraction, the individual patch features are aggregated to form a slide-level representation. This feature aggregation process is facilitated by the MIL aggregator, $f_{a}$. In most cases, $f_{a}$ is trained from scratch (random initialization). This process yields a comprehensive slide-level representation, expressed as $f_{a}(P_{i})$in $\mathbb{R}^{1\times d}$. The primary objective of this step is to conduct the binary classification of slides. Consequently, the final representations are fed into a linear classifier to distinguish between two classes. Each slide is assigned a label $Y_{i}\in\{0,1\}$, where a slide is labeled as $Y_{i}=0$ if it is normal and $Y_{i}=1$ if it contains tumor.
@@ 3.3 Overall training process
Figure 3: Barlow Twins Slide Mixing: This process relies on region features extracted via HIPT feature extractor as inputs, accommodating varying WSI input sizes through zero-padding and masking techniques and the addition of intra-batch mixing strategy during pre-training. “Flip” means from the same batch but with the batch order reversed.
@@
Finally, the most informative WSIs, determined by active learning acquisition functions or random sampling are selected from the unlabeled pool. These selected WSIs are then labeled by pathologists in a weakly-supervised setting, requiring only slide-level labels without fine-grained annotations. Subsequently, they are added to the training data for the next cycle. This process continues until the total labeling budget for WSIs (number of available training WSIs and their labels) is exhausted.